OpenStudy (anonymous):

Show your work. 5 1/3 +(-3 9/18) =

OpenStudy (compassionate):

\[5\frac{ 1 }{3 } + (-3 \frac{ 9 }{ 18 })\] Is that it?

OpenStudy (anonymous):

yes but I can't seem to actually figure it out

OpenStudy (compassionate):

\[\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

so the answer isn't 2 1/45

OpenStudy (compassionate):

No

OpenStudy (anonymous):

ok Thank you

OpenStudy (anonymous):

I believe compassionate is wrong.

OpenStudy (anonymous):

Right now this subtracting mixed fractions.

OpenStudy (anonymous):

So then 1/6 isnt the answer either?

OpenStudy (anonymous):

Lets do this work together. 5 1/3 - 3 9/18 you need to make them into improper fraction do you know how to do this?

OpenStudy (anonymous):

No 1/6 is not the answer I have got.

OpenStudy (anonymous):

no I dont. sorry

OpenStudy (anonymous):

I get 2 1/45

OpenStudy (anonymous):

ok Here I think this will help you a bunch. Tell me what amswer yyou get. http://www.youtube.com/watch?v=yXiG2iv1Vq8

OpenStudy (anonymous):

Answer you*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How is it coming?

OpenStudy (anonymous):

Well I got 1 5/6 if that is what you needed